OpenStudy (anonymous):

\[\int\limits_{c}^{d} \int\limits_{a}^{b} (d^2f)/(dxdy) dxdy\] equal?

OpenStudy (anonymous):

okay the answer is f(b,d)-f(b,c)-f(a,d)+f(a,c)

OpenStudy (anonymous):

right?

OpenStudy (experimentx):

could you please rewrite the question??

OpenStudy (anonymous):

\[\int\limits_{c}^{d} \int\limits_{a}^{b} (d^2f)/(dxdy) dxdy\] equal?

OpenStudy (experimentx):

Is it ?? \[ \int\limits_{c}^{d} \int\limits_{a}^{b} \frac{d^2f}{dxdy} dxdy\] or partials??

OpenStudy (anonymous):

integrate the Partials derivaties

OpenStudy (anonymous):

how the original format?

OpenStudy (experimentx):

\[ \int\limits_{c}^{d} \int\limits_{a}^{b} \frac{\partial^2f}{\partial x\partial y} dxdy \]

OpenStudy (anonymous):

how did you type the dividing and the partical ^^?

OpenStudy (experimentx):

\frac{numerator}{ denominator } \( \partial \) \partial and my super secret technique ... click left, "Show Math as" => "Text Commands" then copy anything

OpenStudy (anonymous):

\partial x \frac{1}{x}

OpenStudy (anonymous):

[\partial x] , [\frac{1}{x}]

OpenStudy (anonymous):

Wow not only did I get beat to the answer but I learned a new formatting technique for this website, thank experimentX, you get a medal from me too :P \[\frac{numerator}{ denominator}\]

OpenStudy (experimentx):

note* This is just a try ... \[ \int\limits_{c}^{d} \int\limits_{a}^{b} \frac{\partial^2f}{\partial x\partial y} dxdy = \int\limits_{c}^{d} \left ( \int\limits_{a}^{b} \frac{\partial }{\partial x} \left ( \frac{\partial f}{\partial y} \right ) dx \right ) dy \] @OsmondF \[ at the beginning

OpenStudy (anonymous):

Test? \[\int\limits\limits_{c}^{d} \int\limits\limits_{a}^{b} \frac{\partial^2f}{\partial x\partial y} dxdy = \int\limits\limits_{c}^{d} \left ( \int\limits\limits_{a}^{b} \frac{\partial }{\partial x} \left ( \frac{\partial f}{\partial y} \right ) dx \right ) dy\]

OpenStudy (experimentx):

Lol ... perfect!!

OpenStudy (anonymous):

Oh! That is nifty, but did you question get answered sufficiently OsmondF?

OpenStudy (anonymous):

where has "show text" ..

OpenStudy (anonymous):

right click?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

\int\limits\limits_{c}^{d} \int\limits\limits_{a}^{b} \frac{\partial^2f}{\partial x\partial y} dxdy = \int\limits\limits_{c}^{d} \left ( \int\limits\limits_{a}^{b} \frac{\partial }{\partial x} \left ( \frac{\partial f}{\partial y} \right ) dx \right ) dy

OpenStudy (anonymous):

""Show Math as" => "Text Commands" then copy anything" <-- there's a pop-up menu for it

OpenStudy (anonymous):

Ah you got to add the \[

OpenStudy (experimentx):

\[ at front

OpenStudy (experimentx):

and \] at back

OpenStudy (anonymous):

\[\int\limits_{c}^{d} \int\limits_{a}^{b} \frac{\partial^2f}{\partial x\partial y} dxdy = \int\limits_{c}^{d} \left ( \int\limits_{a}^{b} \frac{\partial }{\partial x} \left ( \frac{\partial f}{\partial y} \right ) dx \right ) dy

OpenStudy (anonymous):

Don't forget to close the syntax tag too! :D

OpenStudy (anonymous):

\[\int\limits\limits_{c}^{d} \int\limits\limits_{a}^{b} \frac{\partial^2f}{\partial x\partial y} dxdy = \int\limits\limits_{c}^{d} \left ( \int\limits\limits_{a}^{b} \frac{\partial }{\partial x} \left ( \frac{\partial f}{\partial y} \right ) dx \right ) dy\]

OpenStudy (anonymous):

Horray!

OpenStudy (experimentx):

\[ \int\limits\limits_{a}^{b} \frac{\partial }{\partial x} \left ( \frac{\partial f}{\partial y} \right ) dx = \left [ \frac{\partial f}{\partial y}\right ]_a^b = \frac{\partial }{\partial y} f(b, y) - \frac{\partial }{\partial y} f(a, y)\\ \text{ So we have , } \\ \int_c^d \left (\frac{\partial }{\partial y} f(b, y) - \frac{\partial }{\partial y} f(a, y) \right ) dy = \\ \int_c^d \frac{\partial }{\partial y} f(b, y) dy - \int_c^d \frac{\partial }{\partial y} f(a, y) dy = f(b,d) - f(b,c) - (f(a,d) - f(a,c)) \]

OpenStudy (experimentx):

There you go!!